OpenStudy (anonymous):

Simplify. (-5a^3b^6) (-2a^2b0^4 A) 80a^9b^10 B) -80a^9b^10 C) 80a^11b^10 D) -80a^11b^10

OpenStudy (saifoo.khan):

none.

OpenStudy (saifoo.khan):

-5 * -2 = 10

OpenStudy (anonymous):

what you mean none?

OpenStudy (saifoo.khan):

PLease check your question/options are correct.

OpenStudy (anonymous):

YES,WHY? COULD YOU DRAW HOW YOU ANALYSIS?

OpenStudy (jackellyn):

=-5a^3b^6(-2a^2b^4) -5*-2=10 a^3(a^2)=a^5 b^6(b^4)=b^10 =10a^5b^10 which isnt any of the options as @saifoo.khan said

OpenStudy (anonymous):

GET IT

OpenStudy (anonymous):

no lol but ill try on my own